### v1

@@ 3.6 Model Architecture
While the architecture employed throughout MIM research was consistent (a transformer-based encoder and a shallow decoder), there have been some important contributions that further enhance the performance of the pre-training task through architectural modifications [110]. A few attempts have tried to distinguish themselves from the usual ViT-based models, either by using CNNs [111, 112] or by integrating the MIM into the convolution operation [107, 113].
@@
Figure 12: The architecture of MCMAE as presented by [114].
@@
Motivated by leveraging the masked pre-training strategy of autoencoders with convolutional layers, [114] presented an architecture that combines them, as illustrated in Figure 12. The encoder consists of three stages: the first two are convolutional and the last one is transformer-based. First, a mask is sampled to determine which tokens are visible. The mask is then upsampled at the resolutions of the first two stages, in order to be used by masked convolutional blocks. Information from the first two stages is added to the resulting tokens of the third stage, and then, they are linearly projected. Finally, all tokens (predicted and masked) are decoded into the original pixel space. The authors state that the advantage of this method is represented by the multi-scale features learned by the encoder.
@@
SparseMAE [115] offered a novel solution to the problem that small transformers face in not benefiting significantly from MAE pre-training. It did this by concurrently training a full-scale transformer alongside a smaller sparse network, which resides inside the full transformer. This smaller network is tasked with reconstructing the masked patches of input images. Uniquely, SparseMAE independently manages two sets of weights, one for the sparse network and another for the encompassing larger transformer. Despite this separation, both networks aim to achieve the same objective: the accurate reconstruction of the masked image patches. After pre-training, only the sparse network is used for fine-tuning.
Inspired by TokenMoE [116], [110] proposed a pre-training method that is more robust for all downstream tasks. The first step is to obtain the features representations of a pre-trained MAE and cluster them (obtaining the centroids). The architecture of the model is adapted to contain multiple experts (i.e. groups of heads in the transformer layers), each expert being associated with a cluster. The tokens are routed through the expert which the image was assigned to. When applied to a downstream task, the method selects only the most used experts by the dataset for fine-tuning.
In their work, [111] adopted CNNs for masked image modeling. As in the MAE framework, the input is split in non-overlapping patches and some of them are masked. The encoder is composed of sparse convolutional layers, thus preserving the mask pattern intact along the feature maps. The decoder, used for reconstructing the original image, is composed of upsampling blocks that receive the previous layer, as well as the encoder’s features on the same level. The masking regions are replaced with a mask embedding.
@@
[117] proposed to integrate RevCol [118] in the MAE framework. Besides the bottom-up columns, RevCol is extended to include top-down columns as well, thus resembling the encoder-decoder. The network contains reversible connections, helping the model to learn disentangled representations. In this way, the decoder is no longer dropped during fine-tuning, as it contains salient information.
@@ 3.7 Target Features and Objective Function
Several studies have contributed to both the target features and the objective function. The target features are obtained in various ways, mostly by using features deep encoders (e.g. CLIP or  DINO) [119, 120]. The objective function modifications revolve around integrating both reconstruction and contrastive losses.
Figure 13: The approach proposed by [13], which combines MIM with contrastive learning.
MaskCo [13] is a region-level contrastive learning framework. It begins by augmenting images to create two distinct views of the same image, with one view partially masked. The model is then trained using contrastive learning to align the features of the masked regions with those of the corresponding regions in the unmasked view, as depicted in Figure 13.
@@
MaskCLIP [119] combined masked image modeling and the CLIP contrastive loss between images and text in a single framework. Compared with vanilla CLIP, MaskCLIP has an additional loss for reconstructing a masked image and the two losses share the same visual encoder.
[123] presented a two stage distillation framework. In the first stage, the student network distills the task-agnostic knowledge of the teacher, and the authors chose MIM as a proxy task for this goal. Thus, the student learns to align its visible and masked patches representations with those from the teacher. In the second stage, the classic task-oriented knowledge distillation takes place.
[124] reinterpreted MIM via the lens of contrastive learning. They found a formulation showing that the classic MIM approach is equivalent to a setting with Siamese networks, where one network reconstructs the masked tokens and its counterpart focuses on the unmasked tokens. The goal is to closely align the outputs of these models.
To learn better visual representations during MIM pre-training within the medical domain, [90] leveraged an additional modality (text). Two different encoders are used (one for each modality), however, the text encoder is fed the average global vision embedding as well. Their method consists of two reconstruction objectives, one for the image and one for the text, each with its own decoder. Furthermore, the authors employed a third contrastive objective between the fully visible encoded text and a decoded representation (using an additional decoder) of all the image embeddings (masked and unmasked tokens).
@@
[127] improved the MAE pre-training strategy on multimodal data by optimizing the latent encoding space. Some neighboring tokens of the unmasked tokens are sampled as well, in order to compute the reconstruction loss and use the resulting gradient to produce a more explicit latent representation. This is further used to compute the final reconstruction loss. Moreover, an additional contrastive loss is employed to maximize the similarity of the two resulting latent representations of a task, while minimizing it for different tasks.
@@ 3.8 Model Architecture and Objective Function
A number of papers contributed to both the model architecture and the objective function. While the former is the main focus of most works, the latter contribution represents an extension resulting from the architectural modification.
[128] enhanced the standard MAE pipeline by integrating a discriminator for adversarial training. Notably, this discriminator, which shares parameters with the MAE’s encoder, is trained to distinguish between synthesized and real patches. This enhancement is an addition to the existing pipeline, with the typical reconstruction loss still being present in the training process.
@@
In their work, [131] demonstrated how MAE can boost the performance of 3D medical image segmentation. The input volume (a 3D scan) is split into equal subvolumes and these are randomly masked. Then, different views (frontal, horizontal, or longitudinal) are obtained, and a further arbitrary rotation is applied. During pre-training, besides the main reconstruction objective for each view, three more losses are utilized: the rotation angle estimation, a contrastive loss, as well as an additional MSE loss between two different reconstructed views (after being normalized). The architecture of the encoder is based on SwinTransformer. A cross-attention module, which attends to the features between two views, is integrated before the first level of the decoder.
In the work of [132], the audio spectrogram and the images are tokenized. The unmasked tokens are encoded with separate encoders, while also adding a corresponding modality embedding. Then, three separate forward passes are performed through a common encoder: one for each modality embedding, as well as the concatenation of the two. The concatenated tokens, together with the masked tokens, are decoded and the reconstruction loss is applied. Furthermore, a contrastive loss is applied between the averaged-pooled encoding of each modality.
[133] demonstrated that an additional auxiliary supervised classification task helps the MAE pre-training framework. Besides the main reconstruction loss, the authors integrated another branch that takes only a subset of the visible encoded tokens as input, applies an average global pooling operation on them, and finally predicts the class through a multi-layer perceptron. During the fine-tuning part, all tokens are used.
The contribution of [134] is twofold. Firstly, they integrated MAE pre-training for videos by applying a consistency loss between two successive frames that are masked the same, then encoded and decoded with two different networks (one is an exponential moving average of the other). Secondly, they proposed a similar framework, but using sparse convolutional layers instead of a ViT architecture, which results in lower computational costs.
@@ 3.9 Masking Strategy and Target Features
Several studies have impacted both the masking strategy and the target features. In general, these two contributions are jointly proposed, since changing the target features involves a different masking strategy for the new input signal.
[135] presented a method for learning representations useful for object tracking. The method is based on MAE, but the authors use two inputs, one is the search region and the other one is the template. The MAE is trained to reconstruct the search region as it is, and to recreate the template in the position found in the search region.
[136] presented I2P-MAE, a method designed to learn better 3D features by reconstructing masked point clouds. The approach leverages 2D pre-trained models to keep the important point tokens visible while masking. Moreover, the 2D models are used to get the target representations for a semantic reconstruction loss that is applied on the visible tokens, after the decoder.
[137] combined self-supervised knowledge distillation and masked image modeling into a single framework. In the proposed pipeline, the teacher network processes an image from the same class as the student network. The student network processes a masked image, being trained to maximize the similarity between its class token and the teacher’s class token. In addition, the student is trained to distill the knowledge of the most similar tokens of the teacher.
@@
The study of [138] integrated the MAE pipeline into a teacher-student setting for domain adaptive object detection. In this setup, the student network has a dual focus: it learns the detection task using labels generated by the teacher network, and concurrently, it undertakes the reconstruction of missing multi-scale features from the target images. This reconstruction aspect is pivotal, especially when the availability of pseudo-labels from the teacher is limited, as it significantly improves the model’s adaptability to the target domain, ensuring more robust and accurate object detection performance.
@@
[140] utilized the mask-reconstruction strategy in 3D segmentation due to the lack of supervised data, as well as the domain difference between training and testing data. During training, given a pair composed of a 2D image and a 3D point cloud, patches from one modality are masked and the model tries to estimate them using the other modality. A CNN backbone with a lower masking ratio is employed. Having two different datasets (source-labeled and target-unlabeled), the MIM is performed on both datasets, while the supervised task is performed only on the former.
Inspired by the MAE framework, [141] proposed to boost the performance of a ViT model used for object tracking (given an object in a template image, find the same object in the search image) by applying MIM as an additional concurrent task. Both input images are concatenated and passed through the encoder. Besides the main task head, two more decoders are employed. After a high portion of the embeddings is masked, the two sets (each corresponding to an image) are separately fed through one decoder to reconstruct the two frames. The other decoder only receives as input the token embeddings of the search image and reconstructs the template image.
[142] proposed a pre-training method for 3D point clouds by leveraging the corresponding 2D visual representation. The 3D points and their 2D projections are jointly encoded. The resulting encoding is randomly masked and passed through a two-stage decoder. First, there is a shared decoder, then each representation continues through a separate module to reconstruct both input modalities.
Inspired by MAE, [143] modified the MAE framework in order to boost performance when applied to classification downstream tasks in few-shot scenarios. Rather than reconstructing the original pixel space, the authors proposed operating in the latent representation space of a frozen backbone. Two subsets of images, called support and query, are firstly embedded, the latter’s embeddings being masked. Then, the support embeddings are encoded, concatenated with mask tokens, and decoded to estimate the corresponding query embeddings, MSE being the computed reconstruction loss. After being pre-trained with a large labeled dataset, the method is further trained on the smaller dataset (containing few examples per class) in order to learn the global information about each class.
@@ 3.10 Masking Strategy and Model Architecture
A number of studies have influenced both the masking strategy and the employed architecture. The architectural contributions consist of either integrating MIM with different models or adaptations for a specific scenario. These changes lead to different masking strategies that must be adapted. A notable number of papers from this section have multimodal inputs [22, 144, 29, 145].
[146] introduced a set of changes required to be done on the Hierarchical Vision Transformer architectures in order to be compatible with the MAE framework, where the masked tokens are ignored from the input sequence. There are 2 problems when applied directly, one is the window attention with non-overlapped windows and the other is the convolutional and pooling layers. For the first issue, the authors’ solution is to group the tokens from uneven windows with a novel Optimal Grouping algorithm and then apply the mask attention. For the second issue, they opted for using sparse convolutions.
@@
[112] implemented the MAE framework for convolutional networks. One of the changes was to create the masks based on the deep feature maps of the encoder and resize them to the resolution of the input images. The second change was also in the encoder, they used sparse convolutional layers to keep the time performance improvements brought by the masking.
Figure 14: The MAE framework for multiple modalities proposed by [22].
[22] proposed a MAE framework that can be used with multiple modalities. Given a set of modalities, each is tokenized into a common representation form. Rather than using all tokens, only two subsets from each modality are sampled: one that is encoded and the other one that is masked and then reconstructed. A common encoder of all input types is adopted, nevertheless, a modality embedding is added to the tokens. During the cross-attention layers in the decoder, the embeddings corresponding to one modality are masked from the rest while attending all resulting tokens from the encoder. Besides demonstrating good results on downstream tasks, the method shows promising cross-modality generative capabilities.
[144] combined two sources of information (Hematoxylin and Eosin and Immunohistochemical staining images) to detect breast cancer, adopting MAE as the base for their method. Both images are split into patches, some of which will be randomly masked, and the remaining visible patches are fed together through a ViT-based model. The resulting embeddings, together with the learnable mask embeddings, are further processed by two self-attention modules (one specific for each modality), as well as a cross-attention module (that is fed all embeddings). Finally, two separate decoders reconstruct the original images, each receiving the modality-specific embeddings, as well as the inter-modal ones.
@@
PiMAE [29] is a self-supervised framework based on MAE, which learns representations that capture interactions between point clouds and images. Overall, the approach is based on the usual reconstruction objective for each modality. However, in contrast to MAE, the masking strategy in this case is designed to be complementary between the two modalities. In terms of architecture changes, the encoder and decoder include some common blocks between the two modalities, but they also have modality specific layers.
Scale-MAE [148] introduced a pre-training method suitable for scale-dependent domains, specifically in this case they tested on remote sensing data. The method is similar to the MAE framework, but it has some keypoint changes. First, the position encodings that are added to the token embeddings depend on the Earth area covered in the image. Second, the authors changed the decoder to use a three-stage architecture. The first stage is a transformer-based block. The second stage is the upsampling comprising deconvolutional layers. The last stage is called the reconstruction stage and contains Laplacian blocks for reconstructing the high and low frequency features.
In order to leverage the information from multiple input data types to learn richer feature representations, [145] proposed pre-training a network with multiple modalities. Their self-supervised method involved three types of data: an RGB image, depth and semantic segmentation maps. Following the architecture of ViT, the modalities are split into patches and projected into tokens (with separate layers for each). Then, a great portion of the tokens from each modality are masked. Finally, the rest (unmasked) are encoded (using a common encoder), concatenated with the masked tokens, and then a separate decoder for each type is used to reconstruct the corresponding input. During the first stage of the decoder composed of a cross-attention layer, the tokens associated with the respective modality are given as queries, while the keys and values given are from all tokens. Experiments are carried out on downstream tasks for all three types, the pre-training strategy showing competitive results.
[149] made several contributions to adapt the MAE pre-training strategy to their scenario. Firstly, the linear projection layer of the ViT architecture is substituted with CNN layers; however, positional, viewpoint and timestep embeddings are still added. Secondly, they operated on sequences of images that have multiple viewpoints, and thus propose a novel masking strategy: on one hand to fully mask one viewpoint per video frame, on the other hand, the intact frames to have their latent feature maps masked. In order to facilitate the reconstruction task, the encoder is fed with tokens from different viewpoints, as well as from adjacent frames. This pre-training framework allows the authors to train a world model for visual robotic manipulation.
[150] introduced a learnable masking module that facilitates curriculum learning within the MAE framework. Initially the new module creates masks that are easy to reconstruct, the training objective is changed over time to adopt an adversarial role, progressively creating more challenging masks for the MAE to reconstruct. This dynamic adjustment of the training masks enhances the MAE’s ability to handle increasingly complex reconstruction tasks, thereby improving its learning efficiency and robustness.

### v2

@@ 3.6 Model Architecture
While the architecture employed throughout MIM research was consistent (a transformer-based encoder and a shallow decoder), there have been some important contributions that further enhance the performance of the pre-training task through architectural modifications [123]. Some of these advances increase the benefit of MIM pre-training on compact transformers [124, 123], while others integrate multiple modalities within the architecture [125]. Furthermore, a few attempts have tried to distinguish themselves from the usual ViT-based models, either by using CNNs [126, 127, 128] or by integrating MIM into the convolution operation [31, 30].
@@
Figure 16: MCMAE architecture, courtesy of [128] (image licensed under CC BY 4.0). MCMAE falls into the category of reconstruction-based methods illustrated in Figure 2, as it leverages an ${L}^{2}$ loss between the masked patches and the corresponding patches from the input image.
@@
Motivated by leveraging the masked pre-training strategy of autoencoders with convolutional layers, [128] presented an architecture that combines them, as illustrated in Figure 16. The encoder consists of three stages, where the first two are convolutional and the last one is transformer-based. First, a mask is sampled to determine which tokens are visible. The mask is then upsampled at the resolutions of the first two stages, in order to be used by masked convolutional blocks. Information from the first two stages is added to the resulting tokens of the third stage, and then, they are linearly projected. Finally, all tokens (predicted and masked) are decoded into the original pixel space. The authors state that the advantage of this method is represented by the multi-scale features learned by the encoder. Having the same goal of adopting convolutional layers in MIM pipelines, [126] presented a MIM method that leverage CNNs. As in the MAE framework, the input is split in non-overlapping patches and some of them are masked. The encoder is composed of sparse convolutional layers, thus preserving the mask pattern intact along the feature maps. The decoder, used for reconstructing the original image, is composed of upsampling blocks that receive the previous layer, as well as the encoder’s features on the same level. The masking regions are replaced with a mask embedding.
@@
SparseMAE [124] offered a novel solution to the problem that small transformers face in not benefiting significantly from MAE pre-training. It did this by concurrently training a full-scale transformer alongside a smaller sparse network, which resides inside the full transformer. This smaller network is tasked with reconstructing the masked patches of input images. Uniquely, SparseMAE independently manages two sets of weights, one for the sparse network and another for the encompassing larger transformer. Despite this separation, both networks aim to achieve the same objective: the accurate reconstruction of the masked image patches. After pre-training, only the sparse network is used for fine-tuning. A similar approach based on smaller networks is introduced by [123]. Inspired by TokenMoE [129], [123] proposed a pre-training method that is more robust for all downstream tasks. The first step is to obtain the features representations of a pre-trained MAE and cluster them (obtaining the centroids). The architecture of the model is adapted to contain multiple experts (i.e. groups of heads in the transformer layers), each expert being associated with a cluster. The tokens are routed through the expert which the image was assigned to. When applied to a downstream task, the method selects only the most used experts by the dataset for fine-tuning.
@@
[130] proposed to integrate RevCol [131] in the MAE framework. Besides the bottom-up columns, RevCol is extended to include top-down columns as well, thus resembling an encoder-decoder. The network contains reversible connections, helping the model to learn disentangled representations. In this way, the decoder is no longer dropped during fine-tuning, as it contains salient information.
[125] proposed to combine audio and visual modalities at the earliest stage in a multimodal architecture. Aside from separate modality transformer blocks, another block is used to process both the audio and visual tokens, further fusing the resulting information into the other blocks. The experiments demonstrated strong performance on a wide range of downstream tasks, such as visually-guided sound source separation, audio-visual segmentation or classification.
@@ 3.7 Target Features and Objective Function
We next discuss studies that make contributions to both the target features and the objective function. While most approaches use a deep encoder, either frozen [132, 133] or updated online [21, 134, 135, 136, 137] during training, to extract the target features, there are also methods that extract their features directly from the raw input signal [138], eliminating the need for a neural network. In terms of objective functions, most of the works focus on integrating a contrastive loss component [21, 135, 136], but there are also a few methods that leverage a multi-scale reconstruction loss [138].
MaskCo [21] is a region-level contrastive learning framework. It begins by augmenting images to create two distinct views of the same image, with one view partially masked. The model is then trained using contrastive learning to align the features of the masked regions with those of the corresponding regions in the unmasked view. [136] reinterpreted MIM via the lens of contrastive learning. They found a formulation showing that the classic MIM approach is equivalent to a setting with Siamese networks, where one network reconstructs the masked tokens and its counterpart focuses on the unmasked tokens. The goal is to closely align the outputs of these models. MaskCLIP [135] combines masked image modeling and the CLIP contrastive loss between images and text in a single framework. Compared with vanilla CLIP, MaskCLIP has an additional loss for reconstructing a masked image and the two losses share the same visual encoder.
@@
Aside from studying the contrastive objective, additional components added to the loss function were sometimes explored. Interestingly, such studies used multimodal data. To learn better visual representations during MIM pre-training within the medical domain, [137] leveraged an additional modality (text). Two different encoders are used (one for each modality), however, the text encoder is fed with the average global vision embedding as well. The method consists of two reconstruction objectives, one for the image and one for the text, each with its own decoder. Furthermore, the authors employed a third contrastive objective between the fully visible encoded text and a decoded representation (using an additional decoder) of all the image embeddings (masked and unmasked tokens). [139] improved the MAE pre-training strategy on multimodal data by optimizing the latent encoding space. Some neighboring tokens of the unmasked tokens are sampled as well, in order to compute the reconstruction loss and use the resulting gradient to produce a more explicit latent representation. This is further used to compute the final reconstruction loss. Moreover, an additional contrastive loss is employed to maximize the similarity of the two resulting latent representations of a task, while minimizing it for different tasks.
@@
[132] presented a two stage distillation framework. In the first stage, the student network distills the task-agnostic knowledge of the teacher, and the authors chose MIM as a proxy task for this goal. Thus, the student learns to align its visible and masked patches representations with those of the teacher. In the second stage, the classic task-oriented knowledge distillation takes place.
@@ 3.8 Model Architecture and Objective Function
A number of papers contributed to both the model architecture and the objective function. An important subset of these studies made their architectural changes as a result of integrating the new objective [142, 76, 143, 144]. For example, the introduction of adversarial loss by [142] requires an additional discriminator and the multi-scale reconstruction loss used by [76] implies additional adapters to ensure compatibility between features. Another direction of research that requires both architectural and objective function modifications is the integration of multiple modalities [145, 6]. In such scenarios, each modality typically requires its own dedicated modules, leading to architectural adjustments, while an additional loss term is introduced to ensure alignment between the features of the two modalities.
@@
For an improved objective function, some methods were augmented with additional loss components. [144] demonstrated that an additional auxiliary supervised classification task helps the MAE pre-training framework. Besides the main reconstruction loss, the authors integrated another branch that takes only a subset of the visible encoded tokens as input. Further, the branch applies a global average pooling operation on the visible tokens, and finally predicts the class through a multi-layer perceptron. During the fine-tuning stage, all tokens are used. [142] enhanced the standard MAE pipeline by integrating a discriminator for adversarial training. Notably, this discriminator, which shares parameters with the MAE’s encoder, is trained to distinguish between synthesized and real patches. This enhancement is an addition to the existing pipeline, with the typical reconstruction loss still being present in the training process.
@@
[143] demonstrated how MAE can boost the performance of 3D medical image segmentation. The input volume (a 3D scan) is split into equal subvolumes and these are randomly masked. Then, different views (frontal, horizontal, or longitudinal) are obtained, and a further arbitrary rotation is applied. During pre-training, besides the main reconstruction objective for each view, three more losses are utilized: the rotation angle estimation, a contrastive loss, as well as an additional MSE loss between two different reconstructed views (after being normalized). The architecture of the encoder is based on Swin transformer. A cross-attention module, which attends to the features between two views, is integrated before the first level of the decoder.
In the work of [145], the audio spectrogram and the images are tokenized. The unmasked tokens are encoded with separate encoders, while also adding a corresponding modality embedding. Then, three separate forward passes are performed through a common encoder: one for each modality embedding, as well as the concatenation of the two. The concatenated tokens, together with the masked tokens, are decoded and the reconstruction loss is applied. Furthermore, a contrastive loss is applied between the averaged-pooled encoding of each modality. Rather than relying on global information for learning audio-visual features, [6] took a different approach by focusing on a finer-grained level, as well as improving the linkage between the two modalities. Their pipeline involves two more objectives besides the contrastive loss between the embeddings of the two modalities. One additional objective aims to reconstruct the original signal from the unmasked tokens of both input sources, while the second additional objective aims to reconstruct the embeddings using the counterpart tokens and some learnable queries.
The contribution of [147] is twofold. Firstly, they integrated MAE pre-training for videos by applying a consistency loss between two successive frames that are masked in the same manner. The masked frames are then encoded and decoded with two different networks (one is an exponential moving average of the other). Secondly, they proposed an analogous architecture using sparse convolutional layers instead of ViT, which results in lower computational costs.
@@ 3.9 Masking Strategy and Target Features
Several studies have impacted both the masking strategy and the target features. Generally, these studies adjust their target features to reflect changes in masking strategy or in the type of input features that they use [148, 149, 150, 151, 152, 153]. For example, a common scenario is when the model processes pairs of images [148, 154]. In such a scenario, the model employs shared decoders and the target image for reconstruction for at least one of the decoders is altered to reflect the relationship between the images comprising the input pair. Additionally, other studies [149] use the feedback from an auxiliary model to identify the relevant regions for masking and to define the space in which the reconstruction loss is applied.
A couple of papers in this section were applied to object tracking. [148] presented a method for learning representations useful for object tracking. The method is based on MAE, but the authors used two inputs, one is the search region and the other one is the template. The MAE is trained to reconstruct the search region as it is, and to recreate the template in the position found in the search region. Inspired by the MAE framework, [154] proposed to boost the performance of a ViT model used for object tracking (given an object in a template image, find the same object in the search image) by applying MIM as an additional concurrent task. Both input images are concatenated and passed through the encoder. Besides the main task head, two more decoders are employed. After a high portion of the embeddings is masked, the two sets (each corresponding to an image) are separately fed through one decoder to reconstruct the two frames. The other decoder only receives as input the token embeddings of the search image and reconstructs the template image.
Other works focused on implementing MIM for point clouds. On the one hand, [149] presented I2P-MAE, a method designed to learn better 3D features by reconstructing masked point clouds. The approach leverages 2D pre-trained models to keep the important point tokens visible while masking. Moreover, the 2D models are used to get the target representations for a semantic reconstruction loss that is applied on the visible tokens, after the decoder. On the other hand, some works harnessed 2D images in their point-cloud methods. [155] proposed a pre-training method for 3D point clouds by leveraging the corresponding 2D visual representation. The 3D points and their 2D projections are jointly encoded. The resulting encoding is randomly masked and passed through a two-stage decoder. First, there is a shared decoder, then each representation continues through a separate module to reconstruct both input modalities. [152] utilized the mask-reconstruction strategy in 3D segmentation due to the lack of supervised data, as well as the domain difference between training and testing data. During training, given a pair composed of a 2D image and a 3D point cloud, patches from one modality are masked and the model tries to estimate them using the other modality. A CNN backbone with a lower masking ratio is employed. Having two different datasets (source-labeled and target-unlabeled), the MIM is performed on both datasets, while the supervised task is performed only on the former.
@@
The teacher-student framework was utilized in a couple of scenarios. [150] combined self-supervised knowledge distillation and masked image modeling into a single framework. In the proposed pipeline, the teacher network processes an image from the same class as the student network. The student network processes a masked image, being trained to maximize the similarity between its class token and the teacher’s class token. In addition, the student is trained to distill the knowledge of the most similar tokens of the teacher. The study of [151] integrated the MAE pipeline into a teacher-student setting for domain adaptive object detection. In this setup, the student network has a dual focus: it learns the detection task using labels generated by the teacher network, and concurrently, it undertakes the reconstruction of missing multi-scale features from the target images. This reconstruction aspect is pivotal, especially when the availability of pseudo-labels from the teacher is limited, as it significantly improves the model’s adaptability to the target domain, ensuring more robust and accurate object detection performance.
@@
Inspired by MAE, [153] modified the MAE framework in order to boost performance when applied to downstream classification tasks in few-shot scenarios. Rather than reconstructing the original pixel space, the authors proposed to operate in the latent representation space of a frozen backbone. Two subsets of images, called support and query, are firstly embedded, the latter’s embeddings being masked. Then, the support embeddings are encoded, concatenated with mask tokens, and decoded to estimate the corresponding query embeddings, MSE being the computed reconstruction loss. After being pre-trained with a large labeled dataset, the method is further trained on the smaller dataset (containing few examples per class) in order to learn the global information about each class.
@@ 3.10 Masking Strategy and Model Architecture
A number of studies have influenced both the masking strategy and the employed architecture. A notable proportion of these studies have multimodal inputs [26, 157, 158, 159, 160, 161], which typically requires distinct modules for each modality, coupled with carefully designed masking strategies to prevent information leakage between modalities [159, 161]. Additionally, some studies have adapted the MAE framework to be compatible with other popular architectures in computer vision, such as hierarchical transformers [162] and CNNs [127]. Furthermore, there are works that specifically develop architectural modules to directly influence the masking strategy employed during pre-training [51].
@@
While ViT is the main architecture employed in MIM, convolutional layers were still found to be beneficial. [127] implemented the MAE framework for convolutional networks. One of the changes was to create the masks based on the deep feature maps of the encoder and resize them to the resolution of the input images. The second change was also in the encoder, where the authors used sparse convolutional layers to preserve the speed improvements brought by the masking.
[162] introduced a set of changes required by the Hierarchical Vision Transformer architectures in order to be compatible with the MAE framework, where the masked tokens are ignored from the input sequence. There are two problems when applied directly, one is the window attention with non-overlapping windows, and the other is the usage of convolutional and pooling layers. For the first issue, the authors’ solution is to group the tokens from uneven windows with a novel Optimal Grouping algorithm and then apply the mask attention. For the second issue, they opted for using sparse convolutions. [161] made several contributions to adapt the MAE pre-training strategy to their scenario. Firstly, the linear projection layer of the ViT architecture is substituted with CNN layers. However, positional, viewpoint and timestep embeddings are still added. Secondly, the method operates on sequences of images that have multiple viewpoints, requiring a novel masking strategy, on the one hand, to fully mask one viewpoint per video frame, and on the other hand, to mask the latent feature maps of the intact frames. In order to facilitate the reconstruction task, the encoder is fed with tokens from different viewpoints, as well as from adjacent frames. This pre-training framework allows the authors to train a world model for visual robotic manipulation.
Figure 17: The 4M framework for multiple modalities, courtesy of [26] (image available under the ArXiv perpetual license agreement). 4M belongs to the category of reconstruction-based methods shown in Figure 2, as it tries to recover masked discrete tokens using the cross-entropy loss.
Harnessing multiple sources of information during MIM pre-training was demonstrated to bring great benefits, especially for multimodal downstream tasks. [26] proposed an MAE framework that can be used with multiple modalities. Given a set of modalities, each is tokenized into a common representation form. Rather than using all tokens, only two subsets from each modality are sampled: one that is encoded and the other one that is masked and then reconstructed. A joint encoder is employed for all input types. Still, a modality embedding is added to each token to indicate its source. In the cross-attention layers of the decoder, the embeddings corresponding to one modality are masked from the rest, while using all resulting tokens from the encoder. Besides demonstrating good results on downstream tasks, the method shows promising cross-modal generative capabilities. To leverage the information from multiple input data types and learn richer feature representations, [160] proposed to pre-train a network with multiple modalities. Their self-supervised method involved three types of data: an RGB image, a depth map and a semantic segmentation map. Using the ViT architecture, the modalities are split into patches and projected into tokens (with separate layers for each modality). Then, a great proportion of tokens from each modality is masked, while the rest (unmasked) are encoded (using a joint encoder) and concatenated with the masked tokens. A separate decoder for each data type is used to reconstruct the corresponding input. In the first stage of the decoder composed of a cross-attention layer, the tokens associated with the respective modality are given as queries, while the keys and values given are from all tokens. Experiments carried out on downstream tasks for all three data types indicate that the proposed strategy shows competitive results. PiMAE [159] is a self-supervised framework based on MAE, which learns representations that capture interactions between point clouds and images. Overall, the approach is based on the usual reconstruction objective for each modality. However, in contrast to MAE, the masking strategy in this case is designed to be complementary between the two modalities. In terms of architectural changes, the encoder and decoder include some common blocks between the two modalities, but they also have modality specific layers. [157] combined two sources of information (Hematoxylin and Eosin and Immunohistochemical staining images) to detect breast cancer, adopting MAE as the base for their method. Both images are split into patches, some of which are randomly masked, while the remaining visible patches are fed together through a ViT-based model. The resulting embeddings, together with the learnable mask embeddings, are further processed by two self-attention modules (one specific for each modality), as well as a cross-attention module (that processes all embeddings). Finally, two separate decoders reconstruct the original images, each receiving the modality-specific embeddings, as well as the inter-modal ones. More recently, [163] employed the MAE framework to learn robust cross-modal representations. Their primary contribution is the MultiModal 3D Interaction module, which takes as input the concatenation of masked 3D volume features from both video frames and LiDAR sensors. Built upon self-attention layers and feed-forward networks, this module outputs a representation that is then split along the channel dimension and fed into modality-specific decoders to reconstruct the original inputs.
@@
Scale-MAE [164] introduced a pre-training method suitable for scale-dependent domains, specifically testing it on remote sensing data. The method is similar to the MAE framework, but it has some key changes. First, the positional embeddings that are added to the token embeddings depend on the Earth area covered in the image. Second, the authors changed the decoder to use a three-stage architecture. The first stage is a transformer-based block. The second stage performs upsampling with deconvolutional layers. The last stage employs Laplacian blocks to reconstruct the high and low frequency features.
[51] introduced a learnable masking module that facilitates curriculum learning within the MAE framework. Initially, the new module creates masks that are easy to reconstruct. The training objective is changed over time to adopt an adversarial role, progressively creating more challenging masks for the MAE to reconstruct. This dynamic adjustment of the training masks enhances the MAE’s ability to handle increasingly complex reconstruction tasks, thereby improving its learning efficiency and robustness.